The current price of Estelle Corporation stock is $ 23.00. In each of the next two​ years, this stock price will either go up by 16 % or go down by 16 %. The stock pays no dividends. The​ one-year risk-free interest rate is 8.0 % and will remain constant. Using the Binomial​ Model, calculate the price of a​ one-year call option on Estelle stock with a strike price of $ 23.00.

The price of the​ one-year call option is ​$____ (Round to to the nearest​ cent.)
Answer #1
High price=Current price*up move=23*1.16=26.68
Low price=Current price*down move=23*0.84=19.32
Risk neutral probability for up move
q = (e^(risk free rate*time)-D)/(U-D)
=(e^(0.08*1)-0.84)/(1.16-0.84)=0.76027
Call option payoff at high price (payoff H)
=Max(High price-strike price,0)
=Max(26.68-23,0)
=Max(3.68,0)
=3.68
Call option payoff at low price (Payoff L)
=Max(Low price-strike price,0)
=Max(19.32-23,0)
=Max(-3.68,0)
=0
Price of call option = e^(-r*t)*(q*Payoff H+(1-q)*Payoff L)
=e^(-0.08*1)*(0.760272*3.68+(1-0.760272)*0)
=2.58